# Magic dynamics in many-body localized systems

### v2

@@ Abstract
Nonstabilizerness, also known as quantum magic, quantifies the deviation of quantum states from stabilizer states, capturing the complexity necessary for quantum computational advantage. In this study, we investigate the dynamics of quantum magic in disordered many-body localized (MBL) systems using the stabilizer Rényi entropy (SRE). Leveraging a phenomenological description based on the $\ell$-bit model, we analytically and numerically demonstrate that interactions profoundly influence magic spreading, inducing a power-law growth of SRE that markedly contrasts with the rapid saturation observed in ergodic systems. We validate our theoretical predictions through numerical simulations of the disordered transverse-field Ising model, showing excellent agreement across various disorder strengths, system sizes, and initial states. Additionally, we uncover a universal relationship between SRE and entanglement entropy, revealing their common scaling in the MBL regime independent of disorder strength and system size. Our results offer critical insights into the interplay of disorder, interactions, and complexity in quantum many-body systems.
@@ Introduction.
Understanding the generation of magic resources in many-body systems is fundamental for assessing their classical simulability. Generically, quantum many-body systems prepared in an out-of-equilibrium state are expected to follow the eigenstate thermalization hypothesis (ETH[32, 33, 34, 35] and to thermalize reaching an equilibrium state described by appropriate ensembles of statistical mechanics [36, 37, 38, 39]. Thermalization is accompanied by fast, ballistic [40, 41] or sub-ballistic [42, 43], growth of entanglement entropy, and a rapid saturation of nonstabilizerness measures [44, 45] to their maximal values [46]. The process of thermalization slows down in the presence of disorder [47, 48, 49, 50]. Sufficiently strong disorder leads to a phenomenon of many-body localization (MBL[51, 52, 53, 54] which prevents the thermalization [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] at any experimentally relevant time scale [70, 71]. The absence of thermalization starkly affects the dynamics of MBL systems, leading to a logarithmic in time growth of entanglement entropy [72, 73, 74, 75, 76] and the memory of the initial state due to the presence of an emergent set of local integrals of motion, dubbed localized bits or $\ell$-bits [77, 57]. This raises the question: how does nonstabilizerness spread in MBL systems?
@@
Figure 1: Magic spreading in disordered quantum systems. In the non-interacting case (III, red dashed line), $\mathcal{M}_{2}$ is analytically tractable (see [78]). It saturates rapidly to a finite value due to the absence of spin dephasing. In the MBL regime, dephasing induces a power-law growth of $\mathcal{M}_{2}$ toward a saturation value described by Eq. (5). For the initial X-polarized state (I, orange dashed line), magic grows rapidly and saturates to the Haar value. For a generic random product state (II, orange dashed line), the growth is slower and saturates at a lower value, revealing the dependence of magic spreading on the choice of initial state.
@@
In this work, we address this question by investigating the dynamics of nonstabilizerness in strongly disordered spin chains that exhibit MBL. Leveraging the $\ell$-bits, we develop a theoretical framework that describes how nonstabilizerness evolves under MBL dynamics. Specifically, we analyze the stabilizer Rényi entropy (SRE) [28] as a measure of magic, characterizing its growth in both non-interacting and interacting disorder localized systems. In non-interacting localized systems, the SRE remains limited, while in MBL systems, the slow spin dephasing leads to a power-law growth of nonstabilizerness before its eventual saturation. To corroborate our theoretical predictions, we perform numerical simulations on a microscopic MBL model, the disordered transverse-field Ising model (TFIM). We track the evolution of magic across different disorder strengths, initial states, and system sizes, showing remarkable agreement with the $\ell$-bits-based predictions. Additionally, we uncover a universal relationship between the stabilizer Rényi entropy and entanglement entropy, highlighting that their dynamics collapse onto a single curve in the MBL regime, independent of disorder strength and system size.
@@
where $L$ is the number of qubits, $k$ is the Rényi index, and $P$ is a Pauli string that belongs to the Pauli group $\mathcal{P}_{L}$. In particular, $\mathcal{M}_{1}$ is defined by the limit $k\to 1$ in (1), and $\mathcal{M}_{k}\geq 0$, with the equality holding if and only if $|\Psi\rangle$ is a stabilizer state [79, 80]. In our study, we fix $k=2$ and evaluate the SRE using the algorithm of [81], which allowed us to obtain numerically exact results for $L\leq 22$. One advantage of the SRE over many other proposed measures of magic [26] is that it allows an efficient computation even for a large $L$  [82, 83, 84, 85, 86, 87, 88]. Moreover, $\mathcal{M}_{k}(|\Psi\rangle)$ is also experimentally measurable [87, 89, 90, 91, 92].
@@
where $D=2^{L}$ is the Hilbert space dimension, at times scaling logarithmically with system size $L$ [44]. Generic many-body systems exhibit more intricate behavior. Floquet systems behave similarly to random circuits, while for Hamiltonian dynamics, the time required to approach the Haar value scales linearly in $L$, and the SRE may not reach the Haar value [95, 45].
@@
for the initial $|\Psi_{X}^{+}\rangle$ state. In the limit of $t\rightarrow\infty$, this expression yields $\mathcal{M}_{2}\approx L\log_{2}(8/7)$, shown by (III) red dashed line in Fig. 1. For a generic product initial state $|\Psi_{R}\rangle$, our numerical results show that the SRE rapidly saturates to a constant value not bigger than $\mathcal{M}_{2}(|\Psi_{R}\rangle)+L\log_{2}(8/7)$ (see [78] for an analytical derivation for arbitrary initial states)
@@
To assess the accuracy of the $\ell$-bit in capturing the magic dynamics of strongly interacting disordered systems, we analyze a microscopic model expected to exhibit an MBL phase at sufficiently strong disorder, the disordered TFIM, with Hamiltonian
@@
In Fig. 2(a), we show the time evolution of a $Z$ polarized initial state $|\Psi_{Z}\rangle$. For weak disorder, $W\sim 1$, $\mathcal{M}_{2}$ quickly grows towards the Haar value in the weak disorder regime, consistent with the ergodic dynamics results [44, 45]. For strong disorder $W\gtrsim 5$, $|\Psi_{Z}\rangle$ is close to an eigenstate of (6) and, therefore, SRE increases very slowly. In the MBL regime, the behavior of SRE is accurately captured by the phenomenological formula (5) (red dashed line in the Fig. 2(a)), with the saturation value $\mathcal{M}_{2}^{\mathrm{sat}}$ considerably smaller than $\mathcal{M}_{2}^{\mathrm{Haar}}$.
@@
As mentioned before, Pauli strings from $\mathcal{P}_{IZ}$ remain frozen during the dynamics of (3), so $\mathcal{W}_{Z}$ is a constant fixed by the initial state, as shown in Fig. 3(c). Similar behavior is obtained for the TFIM deep in the MBL regime ($W=8$), where $\tau_{i}^{z}\approx Z_{i}$. As illustrated in Fig. 3(d), the choice of $|\Psi_{Y}\rangle$ in the TFIM leads to $W_{Z}\approx 1/D$, in agreement with the completely delocalized case in the $\ell$-bit. On the other hand, the choice of $|\Psi_{Z}\rangle$ leads to a very slow dynamics of $\mathcal{W}_{Z}$, showing that this state is close to an eigenstate of (3) (see the End Matter for further discussion).
@@
We consider $|\Psi_{Y}\rangle$ as our initial state. As shown in Fig. 4 (a), both $\ell$-bit model and TFIM exhibit similar dynamics, with $\mathcal{M}_{2}(\mathcal{S})$ collapsing onto a single master curve without any scaling parameters. For an initial random product state $|\Psi_{R}\rangle$ state the $\ell$-bit model predictions are aligned with the TFIM results when a disorder-dependent rescaling function, $f(W)$, is introduced. In Fig. 4(b) we show $\mathcal{M}_{2}/f(W)$ for $|\Psi_{R}\rangle$, where $f(W)$ is found by minimizing deviations of TFIM results from the $\ell$-bit model predictions. Importantly, $f(W)$ is independent of system size $L$. These results demonstrate a close connection between the SRE and the growth of entanglement in the MBL regime.
@@
In this work, we have explored the dynamics of the nonstabilizerness in disordered many-body systems exhibiting MBL. By developing a theoretical framework founded on an $\ell$-bit phenomenology, we have shown that magic spreading in MBL systems is fundamentally constrained by the slow dynamics characteristic of the MBL regime. Unlike ergodic systems, where SRE rapidly saturates to its maximal value, MBL systems display a much slower power-law relaxation towards the saturation value, which exhibits a strong variability with respect to the choice of initial state. Through numerical simulations of the disordered TFIM, we have verified our phenomenological formulas for the SRE growth and demonstrated a strong connection between the SRE and entanglement entropy growth in the MBL regime. In particular, our findings show that the disorder suppresses the rapid spread of magic resources, and interactions play a crucial role in enabling its slow but sustained growth.
Our results open up several intriguing directions for future research. One promising avenue is investigating the interplay between magic and other forms of ergodicity breaking, such as quantum many-body scars [111, 112, 113] and disorder-free localization [114, 115, 116, 117, 118, 119, 120]. In particular, extending our analysis to gauge theories [121, 122] and constrained quantum systems [123, 124, 125, 126, 127] could shed further light on the role of magic in non-thermalizing quantum dynamics. Moreover, this also motivates further explorations of Clifford-augmented matrix product states [128, 129, 130, 131, 132] in scenarios where continuous Hamiltonians govern the dynamics. Such investigations may provide a broader understanding of the interplay of disorder and magic spreading for quantum error correction and fault-tolerant quantum computing.
@@ Acknowledgments.
Supplementary material: Magic dynamics in many-body localized systems
I Nonstabilizerness of Anderson states
@@
where $\mathbb{E}_{\{h\}}$ denotes the average over different disorder realizations and $\rho$ is the density matrix of the state $|\Psi(t)\rangle$. Since the system is noninteracting, the density matrix can be written as $\rho=\bigotimes_{k=1}^{L}\rho_{k}$, where $\rho_{k}$ is the reduced density matrix of the $k$-th spin. For any initial state $|\Psi\rangle=\bigotimes_{k=1}^{L}(\cos(\theta_{k}/2)|\uparrow\rangle+e^{i\phi_{k}}\sin(\theta_{k}/2)|\downarrow\rangle)$, the reduced density matrix $\rho_{k}$ at time $t$ is given by [Serbyn14]:

### v3

# Nonstabilizerness dynamics in many-body localized systems

@@ Abstract
Nonstabilizerness, also known as “magic”, quantifies the deviation of quantum states from stabilizer states, capturing the complexity necessary for quantum computational advantage. In this study, we investigate the dynamics of nonstabilizerness in disordered many-body localized (MBL) systems using the stabilizer Rényi entropy (SRE). Leveraging a phenomenological description based on the $\ell$-bit model, we analytically and numerically demonstrate that interactions profoundly influence nonstabilizerness spreading, inducing a power-law growth of SRE that markedly contrasts with the rapid saturation observed in ergodic systems. We validate our theoretical predictions through numerical simulations of the disordered transverse-field Ising model, showing excellent agreement across various disorder strengths, system sizes, and initial states. Additionally, we uncover a universal relationship between SRE and entanglement entropy, revealing their common scaling in the MBL regime independent of disorder strength and system size. Our results offer critical insights into the interplay of disorder, interactions, and complexity in quantum many-body systems.
@@ Introduction.
Understanding the generation of magic resources in many-body systems is fundamental for assessing their classical simulability. Generically, quantum many-body systems prepared in an out-of-equilibrium state are expected to follow the eigenstate thermalization hypothesis (ETH) [32, 33, 34, 35] and to thermalize reaching an equilibrium state described by appropriate ensembles of statistical mechanics [36, 37, 38, 39]. Thermalization is accompanied by fast, ballistic [40, 41] or sub-ballistic [42, 43], growth of entanglement entropy, and a rapid saturation of nonstabilizerness measures [44, 45] to their maximal values [46]. The process of thermalization slows down in the presence of disorder [47, 48, 49, 50]. Sufficiently strong disorder leads to a phenomenon of many-body localization (MBL) [51, 52, 53, 54] which prevents the thermalization [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] at any experimentally relevant time scale [70, 71]. The absence of thermalization starkly affects the dynamics of MBL systems, leading to a logarithmic in time growth of entanglement entropy [72, 73, 74, 75, 76] and the memory of the initial state due to the presence of an emergent set of local integrals of motion, dubbed localized bits or $\ell$-bits [77, 57]. This raises the question: how does nonstabilizerness spread in MBL systems?
@@
Figure 1: Nonstabilizerness spreading in disordered quantum systems. In the non-interacting case (III, red dashed line), $\mathcal{M}_{2}$ is analytically tractable (see [78]). It saturates rapidly to a finite value due to the absence of spin dephasing. In the MBL regime, dephasing induces a power-law growth of $\mathcal{M}_{2}$ toward a saturation value described by Eq. (5). For the initial X-polarized state (I, orange dashed line), nonstabilizerness grows rapidly and saturates to the Haar value. For a generic product state (II, orange dashed line), the growth is slower and saturates at a lower value, revealing the dependence of nonstabilizerness spreading on the choice of initial state.
@@
In this work, we address this question by investigating the dynamics of nonstabilizerness in strongly disordered spin chains that exhibit MBL. Leveraging the $\ell$-bits, we develop a theoretical framework that describes how nonstabilizerness evolves under MBL dynamics. Specifically, we analyze the stabilizer Rényi entropy (SRE) [28] as a measure of nonstabilizerness, characterizing its growth in both non-interacting and interacting disorder localized systems. In non-interacting localized systems, the SRE remains limited, while in MBL systems, the slow spin dephasing leads to a power-law growth of nonstabilizerness before its eventual saturation. To corroborate our theoretical predictions, we perform numerical simulations on a microscopic MBL model, the disordered transverse-field Ising model (TFIM). We track the evolution of nonstabilizerness across different disorder strengths, initial states, and system sizes, showing remarkable agreement with the $\ell$-bits-based predictions. Additionally, we uncover a universal relationship between the stabilizer Rényi entropy and entanglement entropy, highlighting that their dynamics collapse onto a single curve in the MBL regime, independent of disorder strength and system size.
@@
where $L$ is the number of qubits, $D=2^{L}$ is the Hilbert space dimension, $k$ is the Rényi index, and $P$ is a Pauli string that belongs to the Pauli group $\mathcal{P}_{L}$. In particular, $\mathcal{M}_{1}$ is defined by the limit $k\to 1$ in (1), and $\mathcal{M}_{k}\geq 0$, with the equality holding if and only if $|\Psi\rangle$ is a stabilizer state [79, 80]. In our study, we fix $k=2$ and evaluate the SRE using the algorithm of [81], which allowed us to obtain numerically exact results for $L\leq 22$. One advantage of the SRE over many other proposed measures of nonstabilizerness [26] is that it allows an efficient computation even for a large $L$ [82, 83, 84, 85, 86, 87, 88]. Moreover, $\mathcal{M}_{k}(|\Psi\rangle)$ is also experimentally measurable [87, 89, 90, 91, 92].
@@
at times scaling logarithmically with system size $L$ [44]. Generic many-body systems exhibit more intricate behavior. Floquet systems behave similarly to random circuits, while for Hamiltonian dynamics, the time required to approach the Haar value scales linearly in $L$, and the SRE may not reach the Haar value [95, 45].
@@
for the initial $|\Psi_{X}^{+}\rangle$ state. In the limit of $t\rightarrow\infty$, this expression yields $\mathcal{M}_{2}\approx L\log_{2}(8/7)$, shown by (III) red dashed line in Fig. 1. For a generic product initial state $|\Psi_{R}\rangle$, our numerical results show that the SRE rapidly saturates to a constant value not bigger than $\mathcal{M}_{2}(|\Psi_{R}\rangle)+L\log_{2}(8/7)$ (see [78] for an analytical derivation for arbitrary initial states).
@@
To assess the accuracy of the $\ell$-bit in capturing the nonstabilizerness dynamics of strongly interacting disordered systems, we analyze a microscopic model expected to exhibit an MBL phase at sufficiently strong disorder, the disordered TFIM, with Hamiltonian
@@
In Fig. 2(a), we show the time evolution of a $Z$ polarized initial state $|\Psi_{Z}\rangle$. For weak disorder, $W\sim 1$, $\mathcal{M}_{2}$ quickly grow towards the Haar value in the weak disorder regime, consistent with the ergodic dynamics results [44, 45]. For strong disorder $W\gtrsim 5$, $|\Psi_{Z}\rangle$ is close to an eigenstate of (6) and, therefore, SRE increases very slowly. In the MBL regime, the behavior of SRE is accurately captured by the phenomenological formula (5) (red dashed line in Fig. 2(a)), with the saturation value $\mathcal{M}_{2}^{\mathrm{sat}}$ considerably smaller than $\mathcal{M}_{2}^{\mathrm{Haar}}$.
@@
As mentioned before, Pauli strings from $\mathcal{P}_{IZ}$ remain frozen during the dynamics of (3), so $\mathcal{W}_{Z}$ is a constant fixed by the initial state, as shown in Fig. 3(c). Similar behavior is obtained for the TFIM deep in the MBL regime ($W=8$), where $\tau_{i}^{z}\approx Z_{i}$. As illustrated in Fig. 3(d), the choice of $|\Psi_{Y}\rangle$ in the TFIM leads to $W_{Z}\approx 1/D$, in agreement with the completely delocalized case in the $\ell$-bit. On the other hand, the choice of $|\Psi_{Z}\rangle$ leads to very slow dynamics of $\mathcal{W}_{Z}$, showing that this state is close to an eigenstate of (3) (see the End Matter for further discussion).
@@
We consider $|\Psi_{Y}\rangle$ as our initial state. As shown in Fig. 4 (a), both $\ell$-bit model and TFIM exhibit similar dynamics, with $\mathcal{M}_{2}(\mathcal{S})$ collapsing onto a single master curve without any scaling parameters. For an initial random product state $|\Psi_{R}\rangle$, the $\ell$-bit model predictions are aligned with the TFIM results when a disorder-dependent rescaling function, $f(W)$, is introduced. In Fig. 4(b) we show $\mathcal{M}_{2}/f(W)$ for $|\Psi_{R}\rangle$, where $f(W)$ is found by minimizing deviations of TFIM results from the $\ell$-bit model predictions. Importantly, $f(W)$ is independent of system size $L$. These results demonstrate a close connection between the SRE and the growth of entanglement in the MBL regime, revealing that the correlations built within the MBL are genuinely quantum, arising from non-Clifford operations. The link between entanglement and nonstabilizerness observed for MBL parallels results in ergodic systems [95, 45, 111], but it is not generic; e.g., scar eigenstates of the PXP model [112] exhibit significant SRE despite limited entanglement (see the End Matter for a more detailed discussion).
@@
In this work, we have explored the dynamics of nonstabilizerness in disordered many-body systems exhibiting MBL. By developing a theoretical framework founded on an $\ell$-bit phenomenology, we have shown that nonstabilizerness spreading in MBL systems is fundamentally constrained by the slow dynamics characteristic of the MBL regime. Unlike ergodic systems, where SRE rapidly saturates to its maximal value, MBL systems display a much slower power-law relaxation towards the saturation value, which exhibits strong variability with respect to the choice of the initial state. Through numerical simulations of the disordered TFIM, we have verified our phenomenological formulas for the SRE growth and demonstrated a strong connection between the SRE and entanglement entropy growth in the MBL regime. In particular, our findings show that the disorder suppresses the rapid spread of magic resources, and interactions play a crucial role in enabling its slow but sustained growth.
Our results open up several intriguing directions for future research. One promising avenue is investigating the interplay between nonstabilizerness and other forms of ergodicity breaking, such as quantum many-body scars [113, 112, 114] and disorder-free localization [115, 116, 117, 118, 119, 120, 121]. In particular, extending our analysis to gauge theories [122, 123] and constrained quantum systems [124, 125, 126, 127, 128] could shed further light on the role of nonstabilizerness in non-thermalizing quantum dynamics. Moreover, this also motivates further explorations of Clifford-augmented matrix product states [129, 130, 131, 132, 133] in scenarios where continuous Hamiltonians govern the dynamics. Such investigations may provide a broader understanding of the interplay of disorder and nonstabilizerness spreading for quantum error correction and fault-tolerant quantum computing.
@@ Acknowledgments.
Data Availability.
- The data that support the findings of this article are openly available [134].
End Matter
In this work, we explored the dynamics of nonstabilizerness, or “magic”, in many-body localized (MBL) systems. Using a combination of analytical and numerical methods, we derived and verified power-law saturation of stabilizer Rényi entropy (SRE) evolution in the MBL regime. Our findings highlight the critical role of interactions and disorder in modulating the growth and saturation of nonstabilizerness, offering clear distinctions between ergodic, Anderson localized, and MBL regimes.
Accuracy of the Power-law Description of the SRE.
In this section, we validate the analytical power-law behavior given by Eq. (5). Specifically, we demonstrate its accuracy by comparing analytical predictions to numerical simulations of SRE dynamics in MBL systems. In order to evaluate the accuracy of (5), we assume $|\Psi_{X}^{+}\rangle$ as the initial state and compute the SRE dynamics for various system sizes $L$, keeping the localization length $\xi=0.5$ (see [78] for more details on the $\ell$-bit Hamiltonian). As illustrated in Fig. 5(a), the SRE exhibits a power-law dependence on time at time-scales where entanglement is relevant, saturating at a size-dependent value (dotted lines). The “bump” observed at early times ($t\sim 1$) is attributed to the initial spin precession that is reminiscent of the non-interacting dynamics. As illustrated in the inset plot of Fig. 5(a), the power-law exponent $\beta^{\prime}$ has a non-trivial system size dependence. This exponent is directly related to the localization length $\xi$ of (3), although its exact relationship remains an outstanding open problem [100, 53].
Figure 5: (a) Nonstabilizerness spread in the $\ell$-bit model for different system sizes, with $|\Psi_{X}^{+}\rangle$ as the initial state. The dashed lines show the analytical solution (Eq. (5) of the main text), which accurately describes the SRE growth after $t\sim 1$. Inset: Dependence of the power-law exponent $\beta^{\prime}$ with the system size for $\xi=0.5$; (b) Time evolution of $\Delta\mathcal{M}_{2}$ for different system sizes $L$. It decays polynomially in $t$ until it eventually saturates, except for $L=18$, for which the Heisenberg time is beyond the considered here. The saturation value exhibits an exponential decay with exponent $\lambda\approx-\ln{2}$, as shown in the inset.
To further analyze the saturation behavior of $\mathcal{M}_{2}$, we introduce the deviation from the Haar value, defined as $\Delta\mathcal{M}_{2}=\mathcal{M}_{2}^{\mathrm{Haar}}-\mathcal{M}_{2}$. Fig. 5(b) illustrates that this saturation value decays exponentially with system size, characterized by the exponent $\lambda\approx-\ln{2}$.
This finite-size dependence observed in the saturation value of the SRE can be attributed primarily to the dynamics of $\mathcal{W}_{Z}$. At sufficiently long times, the spins become completely dephased, making the expectation values of strings involving $X$ and $Y$ gates indistinguishable from those of a random state. However, in finite-size systems, the weight $\mathcal{W}_{Z}$—associated exclusively with identity and $Z$ operators—differs notably. Despite this finite-size discrepancy, we conjecture that both contributions scale as $1/D$ and thus will converge identically in the thermodynamic limit.
Derivation and Dynamics of $\mathcal{W}_{Z}$.
In the main text, we showed that the initial state determines the saturation value of the SRE and, in the strong disorder limit, $\mathcal{M}_{2}$ can be analytically derived. Here, we detail the analytical derivation of the weight of $Z$ gates, $\mathcal{W}_{Z}$, and examine its dynamical behavior in the ergodic and MBL regimes of the TFIM.
The weight of $Z$ gates is defined as:
$\mathcal{W}_{Z}(|\Psi\rangle)=\sum_{P\in\mathcal{P}_{IZ}}\frac{|\langle\Psi|P|\Psi\rangle|^{4}}{D}$ (9)
where $\mathcal{P}_{IZ}$ is the subgroup of the Pauli group $\mathcal{P}_{L}$ consisting exclusively of identity $\hat{I}$ and $\hat{Z}$ operators. Any Pauli operator in this subgroup can be expressed as $P_{w}=\otimes_{k=1}^{L}Z_{k}^{w_{k}}$ with $w\in\{0,1\}^{L}$. For a time-evolved state $|\Psi(t)\rangle=\sum_{u\in\{0,1\}^{L}}c_{u}(t)|u\rangle$, the expectation value of $P_{w}$ is
$\langle\Psi|P_{w}|\Psi\rangle=\sum_{u}|c_{u}(t)|^{2}(-1)^{w\cdot u}|u\rangle$ (10)
Substituting Eq. (10) into Eq. (9), the contribution involving solely $I$ and $Z$ gates is
$\mathcal{W}_{Z}=\frac{1}{2^{L}}\sum_{u,v,k,l}|c_{u}(t)|^{2}|c_{v}(t)|^{2}|c_{k}(t)|^{2}|c_{l}(t)|^{2}\sum_{w}(-1)^{w\cdot s}$ (11)
with $s=u+v+k+l$. The summation over $w$ simplifies to
$\sum_{w\in\{0,1\}^{L}}(-1)^{w\cdot s}=2^{L}\,\delta_{s\bmod 2,0},$ (12)
leading to the constraint $l=u\oplus v\oplus k$. Therefore, $\mathcal{W}_{Z}$ is given by
Figure 6: $\mathcal{W}_{Z}$ dynamics for two different states, $|\Psi_{Y}\rangle$ and $|\Psi_{Z}\rangle$, in the (a) ergodic and (b) localized regimes of the TFIM model. In the simulation, we consider $L=16$ and average the results over, at least, 1000 realizations. In the ergodic regime, the two different states evolve similarly due to the lack of integrals of motion. In the localized regime, however, the weight of $Z$ gates ($\mathcal{W}_{Z}$) depend on the initial state, leading to different saturation values of the SRE.
Since the MBL regime is characterized by an extensive set of integrals of motion, any string from the subgroup $\mathcal{P}_{IZ}$ remains constant throughout the entire evolution. In order to confirm our microscopic model aligns with the $\ell$-bit phenomenology, we must contrast the dynamics of $\mathcal{W}_{Z}$ in both regimes of the TFIM. In Fig. 6(a)), we show that $\mathcal{W}_{Z}$ quickly approaches the value $1/D$ for a random product state in the $Z$ basis, $|\Psi_{Z}\rangle$. For an initial state polarized along the $Y$ basis, $|\Psi_{Y}\rangle$, $\mathcal{W}_{Z}$ remains nearly constant, indicating a delocalized behavior. In contrast, within the MBL regime (Fig. 6(b)), the dynamics significantly differ: for $|\Psi_{Y}\rangle$, $\mathcal{W}_{Z}$ stays near the ergodic value, while for $|\Psi_{Z}\rangle$, it remains almost constant, consistent with the $\ell$-bit phenomenology. Although data suggest a small power-law decay with exponent $\lambda=0.011(2)$ at intermediate times, indicative of potential slow thermalization, extending the analysis to later times shows a diminishing exponent, hinting at eventual saturation at the Heisenberg time.
ETH-MBL crossover-
We now investigate whether $\mathcal{M}_{2}$ can serve as a tool for distinguishing the crossover between ETH and MBL regimes. The SRE has been linked to phase transitions in many ground-state problems [83, 135, 136, 137, 138] and, therefore, one may expect that it is sensitive to the ETH-MBL crossover observed for $L\approx 20$ at $W_{c}\sim 3.5$ [67]. We consider the time evolution of $|\Psi_{Z}\rangle$ and compute $\Delta\mathcal{M}_{2}=\mathcal{M}_{2}^{\mathrm{Haar}}-\mathcal{M}_{2}$ for different size systems $L$ at the longest available time ($t=2\times 10^{4}$), as shown in Fig. 7(a). For weak disorder strength, $\Delta\mathcal{M}_{2}\rightarrow 0$ as $L$ increases, as expected for the ETH regime. On the other hand, $\Delta\mathcal{M}_{2}$ grows linearly with $L$ for strong disorder. The crossover between these two regimes occurs near $W_{c}\approx 3.5$, showing how the SRE distinguishes the ergodic and non-ergodic dynamical regimes.
Relation between entanglement entropy and nonstabilizerness-
One of the central result of this work is the observation that, in the MBL regime, the dynamics of the SRE and the entanglement entropy collapse onto a single universal curve, independent of disorder strength and system size (see Fig. 4).
This strong correlation is highly nontrivial, as the two quantities capture distinct aspects of quantum correlations: entanglement entropy quantifies the amount of bipartite entanglement, while SRE measures the non-Clifford complexity—the genuinely quantum, classically intractable part—of those correlations. Their common scaling thus implies that the slow entanglement growth in MBL systems is accompanied by the gradual accumulation of nonstabilizerness.
Our findings provide a resource-theoretic lens on localization in interacting systems: disorder constrains not only transport and thermalization but also the generation of magic resources. This connection is not universal across all non-ergodic systems. For instance, scarred eigenstates in the PXP model display significant SRE despite limited entanglement [112]. The link between entanglement and nonstabilizerness observed here parallels findings in ergodic systems [95, 45, 111], while its emergence in a non-ergodic regime highlights the genuinely quantum complexity of MBL dynamics.
Figure 7: Nonstabilizerness across the ETH-MBL crossover in TFIM. (a) Deviation from the Haar value $\Delta\mathcal{M}_{2}$ as a function of system size $L$ and for different disorder strengths. (b) $\Delta\mathcal{M}_{2}$ as a function of disorder strength $W$ at the latest time available for the $|\Psi_{Z}\rangle$ distinguishes the ETH and MBL regimes.
@@
Supplementary material: Nonstabilizerness dynamics in many-body localized systems
@@
where $\mathbb{E}_{\{h\}}$ denotes the average over different disorder realizations and $\rho$ is the density matrix of the state $|\Psi(t)\rangle$. Since the system is noninteracting, the density matrix can be written as $\rho=\bigotimes_{k=1}^{L}\rho_{k}$, where $\rho_{k}$ is the reduced density matrix of the $k$-th spin. For any initial state $|\Psi\rangle=\bigotimes_{k=1}^{L}(\cos(\theta_{k}/2)|\uparrow\rangle+e^{i\phi_{k}}\sin(\theta_{k}/2)|\downarrow\rangle)$, the reduced density matrix $\rho_{k}$ at time $t$ is given by [139]: